OpenStudy (anonymous):

What is the distance between the points (5, –6) and (5, 2)? A. 1 B. 3 C. 4 D. 8

OpenStudy (lucaz):

use the pythegorean theorem, the difference from x1 to x2 is a side of a triangle, the difference from y1 to y2 is the other side, so the distance between the two points is the is the hypotenuse (5-5)^2 + (-6-2)^2 = (d)^2 0 + 64 = d^2 d = sqrt(64) = 8